An emf induced by rotating a 1000 turns, 20.0 cm diameter coil in the earth's 5.00×10^-5 T magnetic field. What average emf induced, given the plane of coil is originally perpendicular to the earth's field and is rotated to be parallel to the field in 10.0 ms
1
Expert's answer
2020-03-27T10:52:17-0400

The Faraday's law of electromagnetic induction says


"{\rm emf}=-\frac{\Delta\Phi}{\Delta t}."

The change of magnetic field


"\Delta\Phi=NBA(\cos 90^{\circ}-\cos 0^{\circ})""=-NBA=-\pi r^2NB""=-\pi\times (0.1)^2\times 10^3\times 5\times 10^{-5}=-1.6\times 10^{-3}\rm Wb"

Hence

"{\rm emf}=\frac{1.6\times 10^{-3}}{10\times 10^{-3}}=0.16\:\rm V."